Alvin's father is seven times as old as him. One year ago, he was nine times as old as Alvin. Find Alvin's present age.
Answer by mananth(16946)

let son's age be x
father'e age = 7x
..
1 year ago their ages were
son x-1
father 7x-1
..
7x-1= 9(x-1)
7x-1=9x-9
-2x=-8
x=4 Alvin's age now